### v2

@@ B.3.2 Proofs of auxiliary lemmas
Appendix C Computation
@@
$\min_{\beta\in\mathbf{B},z\in\mathbb{R}^{n}}-\sum_{i=1}^{n}\Gamma_{i}^{\ast}z_{i}$ $\min_{\beta\in\mathbf{B},z\in\mathbb{R}^{n}}-\sum_{i=1}^{n}\Gamma_{i}^{\ast}z_{i}+\overline{\lambda}(\sum_{i=1}^{n}R_{i}^{\ast}z_{i}-n\cdot k)$
$s.t.\ \frac{X_{i}^{\prime}\beta}{C_{i}}\leq z_{i}\leq 1+\frac{X_{i}^{\prime}\beta}{C_{i}},$ $s.t.\ \frac{X_{i}^{\prime}\beta}{C_{i}}\leq z_{i}\leq 1+\frac{X_{i}^{\prime}\beta}{C_{i}},$
@@
$\sum_{i=1}^{n}R_{i}^{\ast}z_{i}<n\cdot k.$ $\sum_{i=1}^{n}R_{i}^{\ast}z_{i}\geq n\cdot k.$
where constants $C_{i}=\sup_{\beta\in B}|X_{i}^{\prime}\beta|$. The additional binary parameters $(z_{1},\dots,z_{n})$ replace the policy functions $g(X_{i})=\mathbf{1}\{\beta^{\prime}X_{i}\geq 0\}$. The sample-analog rule $\widehat{g}_{\text{sample}}$ is the constrained solution on the left hand side. The trade-off rule $\widehat{g}_{\text{tradeoff}}$ is the better solution across constrained and unconstrained policies in (29).
@@
$\mathcal{\widetilde{G}}=\left\{g(x)=\mathbf{1}\{income\cdot 1\{numchild=j\}\leq y_{j}\}:j\in\{0,1,\geq 2\},y_{j}\in\{0,50,100,\dots,500\}\right\}$ (30)
for characteristics $x=(income,\ numchild)$. This grid thus consists of income thresholds every 50% of the federal poverty level, and the thresholds can vary with number of children. I then approximate the infimum over infinite-dimensional $\mathcal{G}$ by the minimum over $\mathcal{\widetilde{G}}$ with estimated covariance i.e. $\underset{g\in\widetilde{\mathcal{G}}}{\min}\ \frac{\widetilde{h}(g)}{\widehat{\Sigma}^{B}(g,g)^{1/2}}$. Here $\widetilde{h}(\cdot)\sim\mathcal{\mathcal{N}}(0,\widehat{\Sigma}^{B})$ is a Gaussian vector indexed by $g\in\widetilde{\mathcal{G}}$, with $\widehat{\Sigma}^{B}$ is sub-matrix of the covariance estimate $\widehat{\Sigma}^{B}(\cdot,\cdot)$ for $g\in\widetilde{\mathcal{G}}$. Based on 10,000 simulation draws I estimate $c_{\alpha}$ to be -2.56 for $\alpha=5\%$. The validity of this approximation is given by the uniform consistency of the covariance estimator under Assumption 3.1.
Second, to select policies in $\hat{\mathcal{G}}_{\alpha}$ is equivalent to invert a test: $\frac{\sqrt{n}\left(\widehat{B}_{n}(g)-k\right)}{\widehat{\Sigma}^{B}(g,g)^{1/2}}\leq c_{\alpha}$. This test inversion can be written as two constraint (note that $c_{\alpha}<0$). The first constraint is linear: $\sum_{i=1}^{n}R_{i}^{\ast}z_{i}\leq n\cdot k$ and the second is $n\cdot\left(\widehat{B}_{n}(g)-k\right)^{2}>c_{\alpha}^{2}\cdot\widehat{\Sigma}^{B}(g,g)$ which translates to a quadratic constraint
$n\left(\frac{\sum_{i=1}^{n}R_{i}^{\ast}z_{i}}{n}-k\right)^{2}>c_{\alpha}^{2}\left(\left(\frac{1}{n}\sum_{i=1}^{n}(R_{i}^{\ast})^{2}z_{i}\right)-\left(\frac{\sum_{i=1}^{n}R_{i}^{\ast}z_{i}}{n}\right)^{2}\right)$
$\Rightarrow$ $\frac{c_{\alpha}^{2}}{n}\sum_{i=1}^{n}(R_{i}^{\ast})^{2}z_{i}+2k\sum_{i=1}^{n}R_{i}^{\ast}z_{i}-\left(\frac{1}{n}+\frac{c_{\alpha}^{2}}{n^{2}}\right)(\sum_{i=1}^{n}R_{i}^{\ast}z_{i})^{2}<nk^{2}$

### v3

@@ B.3.2 Proofs of auxiliary lemmas
Appendix C Additional Theoretical Results and Discussion
Proposition C.1.
Consider the same setting as in Proposition 3.1, except that we have $\operatorname{Pr}_{P}\{C=1\mid X\}\in(0,1)$ almost surely such that the budget function strictly increases for $t\in[\underline{t},\overline{t}]$. Similarly, suppose the budget constraint is at $k=\mathbb{E}_{P}[C\cdot\mathbf{1}\{X\leq\underline{t}\}].$ Then the sample-analog rule $\widehat{g}_{\text{sample}}$ is asymptotically welfare efficient.
Proof.
Proof of Proposition C.1. Since $\operatorname{Pr}_{P}\{C=1|X\}>0$ almost surely, we have $B(t;P)=\mathbb{E}_{P}[C\cdot\mathbf{1}\{X\leq t\}]=\operatorname{Pr}_{P}\{X\leq t\}\cdot\operatorname{Pr}_{P}\{C=1|X\leq t\}$ is strictly increasing in $t$. The constrained optimal threshold is therefore the highest threshold where the constraint is satisfied exactly i.e. $t^{\ast}=\underline{t}.$
Since $X$ is one-dimensional and continuously distributed, the welfare function $W(t;P)$ is continuous in $t$. Since $W(t;P)$ is also strictly increasing in $t$, for all $\epsilon>0$, there exists $t_{\epsilon}<t^{\ast}$ such that
$W(t_{\epsilon};P)=W(t^{\ast};P)-\epsilon.$
Since $C_{i}\cdot\mathbf{1}\{X_{i}\leq t\}$ is bounded for all $t$, the uniform law of large numbers applies and there exists a sequence $b_{n}\rightarrow 0$ such that
$\lim_{n\rightarrow\infty}\operatorname{Pr}_{P^{n}}\left\{\sup_{t}\left|\widehat{B}_{n}(t)-B(t;P)\right|\leq b_{n}\right\}=1.$
Since $t_{\epsilon}<t^{\ast}$ and $B(t;P)$ continuous and strictly increasing, for $n$ large enough (and $b_{n}$ small enough), we must have
$B(t_{\epsilon};P)\leq B(t^{\ast};P)-b_{n}=k-b_{n}<k=B(t^{\ast};P).$
Taken together, we have
$\lim_{n\rightarrow\infty}\operatorname{Pr}_{P^{n}}\left\{\widehat{B}_{n}(t_{\epsilon})\leq k-b_{n}+b_{n}\right\}\geq\lim_{n\rightarrow\infty}\operatorname{Pr}_{P^{n}}\left\{\widehat{B}_{n}(t_{\epsilon})\leq B(t_{\epsilon};P)+b_{n}\right\}=1.$
As in the setting of Proposition 3.1, the policymakers know the welfare function is strictly increasing, and the sample-analog rule solves $\max_{t}\widehat{B}_{n}(t)\text{ subject to }\widehat{B}_{n}(t)\leq k$. However, the solution is not unique because $\widehat{B}_{n}(t)$ is a step function. For the proof, let the sample-analog rule be the largest possible threshold to maximize $\widehat{B}_{n}(t)$:
$\widehat{t}=\max\left\{\arg\max_{t}\{\widehat{B}_{n}(t)\text{ subject to }\widehat{B}_{n}(t)\leq k\}\right\}.$
This implies that $\lim_{n\rightarrow\infty}\operatorname{Pr}_{P^{n}}\left\{t_{\epsilon}\leq\hat{t}\right\}=1$ and since $W(t;P)$ is strictly increasing in $t$, we have
$\lim_{n\rightarrow\infty}\operatorname{Pr}_{P^{n}}\left\{W(\hat{t};P)\geq W(t^{\ast};P)-\epsilon\right\}=\lim_{n\rightarrow\infty}\operatorname{Pr}_{P^{n}}\left\{W(\hat{t};P)\geq W(t_{\epsilon};P)\right\}=1$
which proves the asymptotic welfare efficiency. ∎
Appendix D Additional Empirical Results and Discussion
D.1 Robustness checks
In this section, I present additional empirical results as I vary the significance level $\alpha$ and the penalty level $\bar{\lambda}$ , relative to the benchmark values considered in Section 4.3.2.
I explore how the modified rule $\widehat{g}_{\alpha}$ changes as $\alpha$ varies from 5% (as in the main text) to 50% in Figure D.1 below. Each point on the curve represents the policy chosen by $\widehat{g}_{\alpha}$ for a corresponding statistical significance level $\alpha$. The vertical axis shows the estimated welfare and the horizontal axis shows the estimated budget of the selected policy. As $\alpha$ increases, the statistical test becomes less conservative and the upper confidence band becomes wider. So the feasible region expands and the selected policy achieves higher estimated welfare.
Figure D.1: Frontier of $\widehat{g}_{\alpha}$ across statistical significance levels $\alpha$
I also explore how $\widehat{g}_{\text{tradeoff}}$ changes for a range of values for $\overline{\lambda}$ as illustrated in the Figure D.2 below. Each point on the curve represents the policy chosen by $\widehat{g}_{\text{tradeoff}}$ for a corresponding $\overline{\lambda}$. The vertical axis shows the estimated welfare and the horizontal axis shows the estimated budget achieved by $\widehat{g}_{\text{tradeoff}}$. For example, if each dollar of budget overrun required a repayment of 1.68 times the amount, rather than just 1 as in the main text, $\widehat{g}_{\text{tradeoff}}$ selects a policy that entails no estimated budget violation, aligning with $\widehat{g}_{\text{sample}}$. Because the optimization is discrete in the sample, intermediate values of $\overline{\lambda}$ often result in the same selected policy, leading to gaps in the curve.
Figure D.2: Frontier of $\widehat{g}_{\text{tradeoff}}$ across different penalty levels $\overline{\lambda}$
D.2 Computation
@@
$\min_{\beta\in\mathbf{B},z\in\mathbb{R}^{n}}-\sum_{i=1}^{n}\tau_{i}^{\ast}z_{i}$ $\min_{\beta\in\mathbf{B},z\in\mathbb{R}^{n}}-\sum_{i=1}^{n}\tau_{i}^{\ast}z_{i}+\overline{\lambda}(\sum_{i=1}^{n}C_{i}^{\ast}z_{i}-n\cdot k)$
$\text{ subject to }\frac{X_{i}^{\prime}\beta}{\bar{C}_{i}}\leq z_{i}\leq 1+\frac{X_{i}^{\prime}\beta}{\bar{C}_{i}},$ $\text{ subject to }\frac{X_{i}^{\prime}\beta}{\bar{C}_{i}}\leq z_{i}\leq 1+\frac{X_{i}^{\prime}\beta}{\bar{C}_{i}},$
@@
$\sum_{i=1}^{n}C_{i}^{\ast}z_{i}<n\cdot k.$ $\sum_{i=1}^{n}C_{i}^{\ast}z_{i}\geq n\cdot k.$
where constants $\bar{C}_{i}=\sup_{\beta\in B}|X_{i}^{\prime}\beta|$. The additional binary parameters $(z_{1},\dots,z_{n})$ replace the policy functions $g(X_{i})=\mathbf{1}\{\beta^{\prime}X_{i}\geq 0\}$. The sample-analog rule $\widehat{g}_{\text{sample}}$ is the constrained solution on the left hand side. The trade-off rule $\widehat{g}_{\text{tradeoff}}$ is the better solution across constrained and unconstrained policies in (18).
@@
$\mathcal{\widetilde{G}}=\left\{g(x)=\mathbf{1}\{\text{income}\cdot 1\{\text{numchild}=j\}\leq y_{j}\}:j\in\{0,1,\geq 2\},y_{j}\in\{0,50,100,\dots,500\}\right\}$ (19)
for characteristics $x=(\text{income},\ \text{numchild})$. This grid thus consists of income thresholds every 50% of the federal poverty level, and the thresholds can vary with number of children. I then approximate the infimum over infinite-dimensional $\mathcal{G}$ by the minimum over $\mathcal{\widetilde{G}}$ with estimated covariance i.e. $\underset{g\in\widetilde{\mathcal{G}}}{\min}\ \frac{\widetilde{h}(g)}{\widehat{\Sigma}^{B}(g,g)^{1/2}}$. Here $\widetilde{h}(\cdot)\sim\mathcal{\mathcal{N}}(0,\widehat{\Sigma}^{B})$ is a Gaussian vector indexed by $g\in\widetilde{\mathcal{G}}$, with $\widehat{\Sigma}^{B}$ is sub-matrix of the covariance estimate $\widehat{\Sigma}^{B}(\cdot,\cdot)$ for $g\in\widetilde{\mathcal{G}}$. Based on 10,000 simulation draws I estimate $c_{\alpha}$ to be -2.47 for $\alpha=5\%$. The validity of this approximation is given by the uniform consistency of the covariance estimator under Assumption 3.1.
Second, to select policies in $\hat{\mathcal{G}}_{\alpha}$ is equivalent to invert a test: $\frac{\sqrt{n}\left(\widehat{B}_{n}(g)-k\right)}{\widehat{\Sigma}^{B}(g,g)^{1/2}}\leq c_{\alpha}$. This test inversion can be written as two constraint (note that $c_{\alpha}<0$). The first constraint is linear: $\sum_{i=1}^{n}C_{i}^{\ast}z_{i}\leq n\cdot k$ and the second is $n\cdot\left(\widehat{B}_{n}(g)-k\right)^{2}>c_{\alpha}^{2}\cdot\widehat{\Sigma}^{B}(g,g)$ which translates to a quadratic constraint
$n\left(\frac{\sum_{i=1}^{n}C_{i}^{\ast}z_{i}}{n}-k\right)^{2}>c_{\alpha}^{2}\left(\left(\frac{1}{n}\sum_{i=1}^{n}(C_{i}^{\ast})^{2}z_{i}\right)-\left(\frac{\sum_{i=1}^{n}C_{i}^{\ast}z_{i}}{n}\right)^{2}\right)$
$\Rightarrow$ $\frac{c_{\alpha}^{2}}{n}\sum_{i=1}^{n}(C_{i}^{\ast})^{2}z_{i}+2k\sum_{i=1}^{n}C_{i}^{\ast}z_{i}-\left(\frac{1}{n}+\frac{c_{\alpha}^{2}}{n^{2}}\right)(\sum_{i=1}^{n}C_{i}^{\ast}z_{i})^{2}<nk^{2}$